This is the first year Jenna is playing softball. She has been practicing her batting. On her last swing, the bat made an arc with a radius of 48 inches and swept through 255 degree of a rotation. Assuming the arc is circular, what is the distance the tip of the bat travels to the nearest inch?
1 point
Your answer
1
Expert's answer
2020-04-30T19:37:59-0400

"L=\pi r \frac{\alpha}{180^{\circ}}"

"r=48 in" , "\alpha=255^{\circ}"

Then

"L=\pi 48\frac{255^{\circ}}{180^{\circ}}=68\pi in."